# On the Natural Robustness of Vision-Language Models Against Visual Perception Attacks in Autonomous Driving

### v1

@@ Abstract
Autonomous vehicles (AVs) rely on deep neural networks (DNNs) for critical tasks such as traffic sign recognition (TSR), automated lane centering (ALC), and vehicle detection (VD). However, these models are vulnerable to attacks that can cause misclassifications and compromise safety. Traditional defense mechanisms, including adversarial training, often degrade benign accuracy and fail to generalize against unseen attacks. In this work, we introduce Vehicle Vision Language Models (V2LMs), fine-tuned Vision-Language Models (VLMs) specialized for AV perception. Our findings demonstrate that V2LMs inherently exhibit superior robustness against unseen attacks without requiring adversarial training, maintaining significantly higher accuracy than conventional DNNs under adversarial conditions. We evaluate two deployment strategies: Solo Mode, where individual V2LMs handle specific perception tasks, and Tandem Mode, where a single unified V2LM is fine-tuned for multiple tasks simultaneously. Experimental results reveal that DNNs suffer performance drops of 33%–46% under attacks, whereas V2LMs maintain adversarial accuracy with reductions of less than 8% on average. The Tandem Mode further offers a memory-efficient alternative while achieving comparable robustness to the Solo Mode. Also, we explore the integration of V2LMs as parallel components to AV perception to enhance resilience against adversarial threats. Our results suggest that V2LMs offer a promising path toward more secure and resilient AV perception systems. [^1] The implementation code of this work is available at https://github.com/pedram-mohajer/V2LM
@@ 1 Introduction
Victim Path without V2LMNPC PathVictim Path with V2LM
Figure 1: An attacker casts a shadow on the stop sign, causing the AV with a traditional TSR algorithm to misclassify it and continue driving (red path), leading to a collision. In contrast, with V2LM, the AV remains robust and correctly stops (blue path) despite the attack.
Autonomous vehicles (AVs) are built on automated driving systems (ADS), consisting of modular subsystems including perception, planning, and control. The perception module—significantly enhanced by advancements in deep neural network (DNN) models for tasks such as image classification and object detection [59]—reads and processes data from sensors such as cameras and LiDAR; this information is then sent to the planning module, which makes decisions on navigation and maneuvers, and subsequently to the control module, executing these decisions to control the vehicle’s movement [49].
@@
Various defense methods have been proposed to mitigate such attacks: defensive distillation reduces model sensitivity but struggles to generalize across diverse perturbations [41]; input transformations can suppress adversarial noise but often degrade clean data quality, lowering accuracy [20]; and provable defenses, though theoretically robust, are computationally expensive and difficult to scale [60]. Among these, adversarial training is the most widely adopted defense. Although it improves robustness against specific attacks, it often reduces generalization and standard accuracy on clean data [43], as shown in 2(a), which is especially problematic in AV contexts where high performance under normal conditions is essential for safe and reliable operation [58].
BenignAdversarial000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111AccuracyNVILAQwen-VLYolov5-cls
(a) After Adversarial Training. Adversarial training improves robustness but reduces benign accuracy, especially for YOLOv5-cls.
BenignAdversarial000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Accuracy
(b) Before Adversarial Training. Fine-tuned VLMs (NVILA, Qwen-VL) exhibit strong robustness even without adversarial training.
Figure 2: Accuracy comparison of NVILA, Qwen-VL, and YOLOv5-cls on benign and adversarial inputs before and after adversarial training. VLMs maintain high robustness with minimal trade-off, while traditional DNNs like YOLOv5-cls show greater vulnerability and poorer generalization.
@@
To address these limitations and achieve both robust and generalizable models for AV perception tasks, this work proposes a novel finding: fine-tuned VLMs inherently exhibit robustness against unseen adversarial attacks on AV perception systems even without adversarial training, significantly outperforming DNNs. As shown in Figure 2(b), VLMs such as Qwen-VL [4] and NVILA [32] demonstrate strong resilience, achieving 82.93% and 98.04% accuracy respectively against adversarial examples without adversarial training. In contrast, the YOLOv5-cls [26, 44] model suffers a drastic drop from 94.88% accuracy on benign inputs to just 55.82% on adversarial ones. Based on this observation, we introduce fine-tuned VLMs for the first time as a robust AV perception module, naming them Vehicle Vision Language Models (V2LMs) and comprehensively evaluating them in the AV context.
@@
This paper evaluates six VLMs, namely LLaVA-7B, LLaVA-13B-LoRA [30], MoE-LLaVA [28], MobileVLM [15], Qwen-VL-7B, and NVILA-8B for their potential as auxiliary in-vehicle components to enhance perception performance in the presence of attacks targeting perception algorithms. The first four models are based on LLaMA [56], while the last is based on Qwen [3], selected for its state-of-the-art performance in multimodal tasks and its efficiency in offline operation—an essential feature for our application. In addition to these VLMs, the study also evaluates task-specific DNN models, including YOLOv5-cls [26, 44] for Traffic Sign Recognition (TSR), CLRerNet [21] for Automated Lane Centering (ALC), and YOLOv5-dt for Vehicle Detection (VD), as strong baselines under both benign and adversarial conditions.
@@
Then, the effectiveness and resilience of both DNN models and V2LMs are tested using unseen adversarial examples (AEs) against three distinct attacks targeting AV perception algorithms (RQ2): (1) Robust and Accurate UV-map-based Camouflage attack (RAUCA) to deceive VD algorithms [70], (2) a physical-world adversarial attack known as the Dirty Road Patch (DRP-Attack) to compromise DNN-based automated lane centering (ALC) models [48], and (3) shadows cast on traffic signs to attack TSR algorithms [69]. As shown in Figure 1, the V2LM improves robustness against an adversarial attack on a traffic sign, helping the AV avoid a collision.
@@
  • This work proposes a novel finding that fine-tuned Vision Language Models (VLMs) inherently exhibit superior robustness against unseen adversarial attacks compared to task-specific DNNs. We highlight a critical limitation of traditional defense methods, such as adversarial training: while adversarial training aims to improve robustness, it significantly degrades the benign accuracy of traditional DNNs while providing only limited improvements against adversarial examples. In contrast, VLM-based models achieve strong adversarial robustness while maintaining high benign accuracy.
  • We introduce Vehicle Vision Language Models (V2LMs), fine-tuned VLMs specifically for AV perception tasks: TSR, ALC, and VD. We further propose two deployment strategies: Solo Mode, where separate VLM are fine-tuned individually for each task, and Tandem Mode, employing a single unified V2LM across multiple perception tasks in same time.
  • We conduct comprehensive experiments to evaluate the robustness of traditional DNN models and V2LMs under adversarial conditions. Traditional DNN models experience performance drops of 33%–46% under attacks, whereas V2LMs achieve significantly smaller reductions (less than 8% on average), consistently maintaining high adversarial accuracy without relying on additional defense mechanisms.
@@ 3.3 Vehicle Vision Language Model (V2LM)
Deploying V2LMs to enhance robustness against AEs in AVs raises important considerations for their implementation. Given the diverse range of tasks that AVs must perform, it is crucial to determine the best strategy for utilizing them. One approach involves using a separate V2LM for each specific AV task to improve robustness within each module, ensuring specialized and precise detection capabilities. This design, referred to as Solo Mode, is illustrated in 3(a), where individual V2LMs are dedicated to tasks such as TSR, ALC, and VD. The formal representation of this design is:
@@
Alternatively, a single V2LM can handle all AV tasks in a tandem approach, providing a unified method for improving robustness across multiple modules, as depicted in 3(b). In this design, multiple image-query pairs—each corresponding to a different task—are combined into a single input, where images are concatenated using a separator, and queries are merged in the same order. The model processes this structured batch simultaneously within one forward pass, extracting task-specific outputs for each image-query pair independently while treating the collection as a cohesive input during execution:
@@ 4.1 Experimental Setup
Accuracy F1-Score Precision Recall Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 95.93% 97.38% 1.45% 94.62% 91.22% -3.40% 94.94% 91.08% -3.86% 94.31% 91.37% -2.94% LLaVA-13B 97.15% 98.14% 0.99% 96.09% 93.21% -2.88% 96.25% 92.35% -3.90% 95.94% 94.10% -1.84% MoE-LLaVA 95.24% 96.91% 1.67% 94.73% 95.78% 1.05% 94.51% 95.26% 0.75% 94.96% 96.31% 1.35% MobileVLM 88.19% 90.34% 2.15% 87.68% 90.18% 2.50% 86.17% 90.73% 4.56% 89.25% 89.65% 0.40% Qwen-7B 95.80% 94.13% -1.67% 95.78% 94.27% -1.51% 95.96% 94.63% -1.33% 95.64% 93.88% -1.76% NVILA-8B 99.04% 99.14% 0.1% 99.10% 99.20% 0.1% 99.11% 99.19% 0.08% 99.13% 99.14% 0.01% ALC LLaVA-7B 98.11% 95.41% -2.70% 98.05% 95.24% -2.81% 98.20% 95.21% -2.99% 97.91% 95.29% -2.62% LLaVA-13B 99.30% 98.31% -0.99% 99.46% 97.85% -1.61% 99.46% 98.32% -1.14% 99.47% 97.40% -2.07% MoE-LLaVA 96.86% 92.53% -4.33% 96.23% 91.83% -4.40% 97.63% 92.37% -5.26% 96.85% 91.30% -5.55% MobileVLM 84.41% 86.61% 2.20% 87.68% 89.53% 1.85% 86.17% 88.93% 2.76% 89.26% 90.10% 0.84% Qwen-7B 93.75% 93.91% 0.16% 93.83% 93.76% -0.07% 94.39% 94.48% 0.09% 93.27% 93.05% -0.22% NVILA-8B 99.51% 99.41% -0.1% 99.45% 99.52% 0.07% 99.34% 99.44% 0.1% 99.55% 99.61% -0.11% VD LLaVA-7B 95.84% 91.90% -3.94% 94.56% 92.82% -1.74% 95.12% 94.71% -0.41% 94.01% 91.21% -2.80% LLaVA-13B 97.05% 98.96% 1.91% 96.76% 95.97% -0.79% 97.33% 97.88% 0.55% 96.21% 94.14% -2.07% MoE-LLaVA 95.52% 93.42% -2.10% 94.78% 92.30% -2.48% 95.31% 93.83% -1.48% 94.27% 90.82% -3.45% MobileVLM 86.61% 88.01% 1.40% 88.42% 89.38% 0.96% 86.76% 87.52% 0.76% 90.15% 91.34% 1.19% Qwen-7B 96.02% 95.01% -1.01% 96.55% 96.17% -0.38% 97.76% 97.35% -0.41% 94.91% 95.01% 0.10% NVILA-8B 99.93% 99.95% 0.02% 95.90% 99.84% -0.06% 99.81% 99.79% -0.02% 99.88% 99.91% 0.1%
@@ 4.2 RQ1: Fine-Tuning Increases Detection Performance
Figure 6: IoU values before and after fine-tuning.
@@ 4.3 RQ2: V2LMs Demonstrate Robustness under Attacks
Accuracy F1-Score Precision Recall Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 80.12% 86.51% 6.39% 86.44% 85.89% -0.55% 85.24% 86.12% 0.88% 87.69% 85.67% 2.02% LLaVA-13B 86.53% 89.01% 2.48% 86.08% 88.95% 2.87% 85.13% 89.31% 4.18% 87.06% 88.60% 1.54% MoE-LLaVA 80.03% 81.52% 1.49% 80.75% 82.34% 1.59% 82.13% 83.26% 1.13% 79.43% 81.45% 2.02% MobileVLM 79.57% 79.13% -0.44% 77.55% 79.88% 2.33% 78.02% 80.75% 2.73% 77.10% 79.04% 1.94% Qwen-7B 82.83% 81.66% -1.17% 83.16% 83.43% 0.27% 85.73% 86.52% 0.79% 80.57% 80.77% 0.20% NVILA-8B 86.04% X-X87.15% X-X90.28% X-X86.09% X-XALC LLaVA-7B 86.07% 84.83% -1.24% 87.01% 85.38% -1.63% 86.68% 84.65% -2.03% 87.53% 86.13% -1.40% LLaVA-13B 90.05% 86.69% -3.36% 90.69% 87.38% -3.31% 89.67% 86.34% -3.33% 91.75% 88.46% -3.29% MoE-LLaVA 82.54% 82.26% -0.28% 83.27% 83.06% -0.21% 84.46% 82.49% -1.97% 82.13% 83.65% 1.52% MobileVLM 78.21% 78.23% 0.02% 80.92% 81.21% 0.29% 83.16% 81.96% -1.20% 78.80% 80.49% 1.69% Qwen-7B 88.81% 86.57% -2.24% 89.48% 85.48% -4% 91.06% 89.61% -1.45% 88.01% 86.57% -1.44% NVILA-8B 99.25% X-X99.13% X-X99.01% X-X99.00% X-XVD LLaVA-7B 89.23% 88.02% -1.21% 87.84% 86.59% -1.25% 89.06% 87.89% -1.17% 86.67% 85.33% -1.34% LLaVA-13B 91.82% 90.09% -1.73% 90.72% 90.41% -0.31% 90.10% 91.03% 0.93% 91.35% 89.80% -1.55% MoE-LLaVA 88.46% 86.54% -1.92% 87.16% 86.21% -0.95% 88.28% 86.77% -1.51% 86.07% 85.67% -0.40% MobileVLM 80.13% 80.15% 0.02% 82.88% 82.97% 0.09% 84.14% 83.81% -0.33% 81.67% 82.16% 0.49% Qwen-7B 89.05% 89.06% 0.01% 90.15% 90.57% 0.42% 92.83% 92.77% -0.06% 87.60% 88.46% 0.86% NVILA-8B 99.81% X-X99.82% X-X99.87% X-X99.81% X-X%

### v2

@@ Abstract
Autonomous vehicles rely on deep neural networks for critical tasks such as traffic sign recognition, automated lane centering, and vehicle detection. However, these models are vulnerable to attacks that can cause misclassifications and compromise safety. Defense mechanisms, including adversarial training, often degrade benign accuracy and fail to generalize against unseen attacks. In this work, we introduce Vehicle Vision Language Models (V2LMs), fine-tuned Vision-Language Models (VLMs) specialized for AV perception. Our findings demonstrate that V2LMs inherently exhibit superior robustness against unseen attacks without requiring adversarial training, maintaining significantly higher accuracy than conventional DNNs under adversarial conditions. We evaluate two deployment strategies: Solo Mode, where individual V2LMs handle specific perception tasks, and Tandem Mode, where a single unified V2LM is fine-tuned for multiple tasks simultaneously. Experimental results reveal that DNNs suffer performance drops of 33%–46% under attacks, whereas V2LMs maintain adversarial accuracy with reductions of less than 8% on average. The Tandem Mode further offers a memory-efficient alternative while achieving comparable robustness to the Solo Mode. Also, we explore the integration of V2LMs as parallel components to AV perception to enhance resilience against adversarial threats. Our results suggest that V2LMs offer a promising path toward more secure and resilient AV perception systems. [^1] The implementation code of this work is available at https://github.com/pedram-mohajer/V2LM
@@ 1 Introduction
Autonomous vehicles (AVs) are built on automated driving systems (ADS), consisting of modular subsystems including perception, planning, and control. The perception module—significantly enhanced by advancements in deep neural network (DNN) models for tasks such as image classification and object detection [62, 40]—reads and processes data from sensors such as cameras and LiDAR [41, 44]; this information is then sent to the planning module, which makes decisions on navigation and maneuvers, and subsequently to the control module, executing these decisions to control the vehicle’s movement [52].
@@
Various defense methods have been proposed to mitigate such attacks: defensive distillation reduces model sensitivity but struggles to generalize across diverse perturbations [43]; input transformations can suppress adversarial noise but often degrade clean data quality, lowering accuracy [20]; and provable defenses, though theoretically robust, are computationally expensive and difficult to scale [63]. Among these, adversarial training is the most widely adopted defense [36].
1(a) shows that adversarial training (AT) reduces YOLOv5-cls’s accuracy on benign inputs by over 20%, while NVILA-8B and Qwen-VL-7B remain stable, with changes under 3%—suggesting that, unlike conventional models, VLMs are less sensitive to the trade-offs typically introduced by AT [46]. 1(b) shows that both VLMs also achieve high accuracy on adversarial inputs, with NVILA-8B staying near 98% and Qwen-VL-7B near 85%. In contrast, YOLOv5-cls starts at 55.8% under attack and gains only 13.2% with AT [61].
@@
To address these limitations and achieve both robust and generalizable models for AV perception tasks, this work proposes a novel finding: fine-tuned VLMs inherently exhibit robustness against unseen adversarial attacks on AV perception systems even without adversarial training, significantly outperforming DNNs. As shown in 1(b), VLMs such as Qwen-VL [4] and NVILA [32] demonstrate strong resilience, achieving 82.93% and 98.04% accuracy respectively against adversarial examples without adversarial training. In contrast, the YOLOv5-cls [26, 47] model suffers a drastic drop from 94.88% accuracy on benign inputs to just 55.82% on adversarial ones. Based on this observation, we introduce fine-tuned VLMs for the first time as a robust AV perception module, naming them Vehicle Vision Language Models (V2LMs) and comprehensively evaluating them in the AV context.
@@
This paper evaluates six VLMs, namely LLaVA-7B, LLaVA-13B-LoRA [30], MoE-LLaVA [28], MobileVLM [15], Qwen-VL-7B, and NVILA-8B for their potential as auxiliary in-vehicle components to enhance perception performance in the presence of attacks targeting perception algorithms. The first four models are based on LLaMA [59], while the last two—Qwen-VL-7B and NVILA-8B—are based on Qwen [3], selected for their state-of-the-art performance in multimodal tasks and their efficiency in offline operation—an essential feature for our application. In addition to these VLMs, the study also evaluates task-specific DNN models, including YOLOv5-cls [26] for Traffic Sign Recognition (TSR), CLRerNet [21] for Automated Lane Centering (ALC), and YOLOv5-dt [47] for Vehicle Detection (VD), as strong baselines under both benign and adversarial conditions.
Before ATAfter AT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111AccuracyNVILAQwen-VLYOLOv5-cls
(a) Benign
Before ATAfter AT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Accuracy
(b) Adversarial
Figure 1: Impact of adversarial training (AT) on NVILA, Qwen-VL, and YOLOv5-cls. (a) Benign inputs: AT reduces clean accuracy for YOLOv5-cls while VLMs remain high. (b) Adversarial inputs: VLMs (NVILA/Qwen-VL) exhibit strong robustness even before adversarial training; YOLOv5-cls still lags.
@@
Then, the effectiveness and resilience of both DNN models and V2LMs are tested using unseen adversarial examples (AEs) against three distinct attacks targeting AV perception algorithms (RQ2): (1) Robust and Accurate UV-map-based Camouflage attack (RAUCA) to deceive VD algorithms [73], (2) a physical-world adversarial attack known as the Dirty Road Patch (DRP-Attack) to compromise DNN-based automated lane centering (ALC) models [51], and (3) shadows cast on traffic signs to attack TSR algorithms [72].
@@
  • This work proposes a novel finding that fine-tuned Vision Language Models (VLMs) inherently exhibit superior robustness against unseen adversarial attacks compared to task-specific DNNs. We highlight a critical limitation of defense methods, such as adversarial training: while adversarial training aims to improve robustness, it significantly degrades the benign accuracy of DNNs while providing only limited improvements against adversarial examples. In contrast, VLM-based models achieve strong adversarial robustness while maintaining high benign accuracy.
  • We introduce Vehicle Vision Language Models (V2LMs)—fine-tuned VLMs specifically for AV perception tasks: TSR, ALC, and VD. We further propose two deployment strategies: Solo Mode, where separate VLM are fine-tuned individually for each task, and Tandem Mode, employing a single unified V2LM across multiple perception tasks in same time.
  • We conduct comprehensive experiments to evaluate the robustness of DNN models and V2LMs under adversarial conditions. DNN models experience performance drops of 33%–46% under attacks, whereas V2LMs achieve significantly smaller reductions (less than 8% on average), consistently maintaining high adversarial accuracy without relying on additional defense mechanisms.
@@ 3.3 Vehicle Vision Language Model (V2LM)
Deploying V2LMs to enhance robustness against AEs in AVs raises important considerations for their implementation. Given the diverse range of tasks that AVs must perform, it is crucial to determine the best strategy for utilizing them. One approach involves using a separate V2LM for each specific AV task to improve robustness within each module, ensuring specialized and precise detection capabilities. This design, referred to as Solo Mode, is illustrated in 2(a), where individual V2LMs are dedicated to tasks such as TSR, ALC, and VD. The formal representation of this design is:
@@
Alternatively, a single V2LM can handle all AV tasks in a tandem approach, providing a unified method for improving robustness across multiple modules, as depicted in 2(b). In this design, multiple image-query pairs—each corresponding to a different task—are combined into a single input, where images are concatenated using a separator, and queries are merged in the same order. The model processes this structured batch simultaneously within one forward pass, extracting task-specific outputs for each image-query pair independently while treating the collection as a cohesive input during execution:
@@ 4.1 Experimental Setup
Accuracy F1-Score Precision Recall Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 95.93% 97.38% 1.45% 94.62% 91.22% -3.40% 94.94% 91.08% -3.86% 94.31% 91.37% -2.94% LLaVA-13B 97.15% 98.14% 0.99% 96.09% 93.21% -2.88% 96.25% 92.35% -3.90% 95.94% 94.10% -1.84% MoE-LLaVA 95.24% 96.91% 1.67% 94.73% 95.78% 1.05% 94.51% 95.26% 0.75% 94.96% 96.31% 1.35% MobileVLM 88.19% 90.34% 2.15% 87.68% 90.18% 2.50% 86.17% 90.73% 4.56% 89.25% 89.65% 0.40% Qwen-7B 95.80% 94.13% -1.67% 95.78% 94.27% -1.51% 95.96% 94.63% -1.33% 95.64% 93.88% -1.76% NVILA-8B 99.04% 99.14% 0.1% 99.10% 99.20% 0.1% 99.11% 99.19% 0.08% 99.13% 99.14% 0.01% ALC LLaVA-7B 98.11% 95.41% -2.70% 98.05% 95.24% -2.81% 98.20% 95.21% -2.99% 97.91% 95.29% -2.62% LLaVA-13B 99.30% 98.31% -0.99% 99.46% 97.85% -1.61% 99.46% 98.32% -1.14% 99.47% 97.40% -2.07% MoE-LLaVA 96.86% 92.53% -4.33% 96.23% 91.83% -4.40% 97.63% 92.37% -5.26% 96.85% 91.30% -5.55% MobileVLM 84.41% 86.61% 2.20% 87.68% 89.53% 1.85% 86.17% 88.93% 2.76% 89.26% 90.10% 0.84% Qwen-7B 93.75% 93.91% 0.16% 93.83% 93.76% -0.07% 94.39% 94.48% 0.09% 93.27% 93.05% -0.22% NVILA-8B 99.51% 99.41% -0.1% 99.45% 99.52% 0.07% 99.34% 99.44% 0.1% 99.55% 99.61% 0.06% VD LLaVA-7B 95.84% 91.90% -3.94% 94.56% 92.82% -1.74% 95.12% 94.71% -0.41% 94.01% 91.21% -2.80% LLaVA-13B 97.05% 98.96% 1.91% 96.76% 95.97% -0.79% 97.33% 97.88% 0.55% 96.21% 94.14% -2.07% MoE-LLaVA 95.52% 93.42% -2.10% 94.78% 92.30% -2.48% 95.31% 93.83% -1.48% 94.27% 90.82% -3.45% MobileVLM 86.61% 88.01% 1.40% 88.42% 89.38% 0.96% 86.76% 87.52% 0.76% 90.15% 91.34% 1.19% Qwen-7B 96.02% 95.01% -1.01% 96.55% 96.17% -0.38% 97.76% 97.35% -0.41% 94.91% 95.01% 0.10% NVILA-8B 99.93% 99.95% 0.02% 99.83% 99.84% 0.01% 99.81% 99.79% -0.02% 99.88% 99.91% 0.03%
@@ 4.2 RQ1: Fine-Tuning Increases Detection Performance
Figure 5: IoU values before and after fine‐tuning.
@@ 4.3 RQ2: V2LMs Demonstrate Robustness under Attacks
Accuracy F1-Score Precision Recall Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 80.12% 86.51% 6.39% 86.44% 85.89% -0.55% 85.24% 86.12% 0.88% 87.69% 85.67% -2.02% LLaVA-13B 86.53% 89.01% 2.48% 86.08% 88.95% 2.87% 85.13% 89.31% 4.18% 87.06% 88.60% 1.54% MoE-LLaVA 80.03% 81.52% 1.49% 80.75% 82.34% 1.59% 82.13% 83.26% 1.13% 79.43% 81.45% 2.02% MobileVLM 79.57% 79.13% -0.44% 77.55% 79.88% 2.33% 78.02% 80.75% 2.73% 77.10% 79.04% 1.94% Qwen-7B 82.83% 81.66% -1.17% 83.16% 83.43% 0.27% 85.73% 86.52% 0.79% 80.57% 80.77% 0.20% NVILA-8B 86.04% 91.12% 5.08% 87.15% 91.10% 3.95% 90.28% 91.15% 0.87% 86.09% 91.10% 5.01% ALC LLaVA-7B 86.07% 84.83% -1.24% 87.01% 85.38% -1.63% 86.68% 84.65% -2.03% 87.53% 86.13% -1.40% LLaVA-13B 90.05% 86.69% -3.36% 90.69% 87.38% -3.31% 89.67% 86.34% -3.33% 91.75% 88.46% -3.29% MoE-LLaVA 82.54% 82.26% -0.28% 83.27% 83.06% -0.21% 84.46% 82.49% -1.97% 82.13% 83.65% 1.52% MobileVLM 78.21% 78.23% 0.02% 80.92% 81.21% 0.29% 83.16% 81.96% -1.20% 78.80% 80.49% 1.69% Qwen-7B 88.81% 86.57% -2.24% 89.48% 85.48% -4% 91.06% 89.61% -1.45% 88.01% 86.57% -1.44% NVILA-8B 99.25% 99.14% –0.11% 99.13% 99.10% –0.03% 99.01% 99.11% +0.10% 99.00% 99.10% +0.10% VD LLaVA-7B 89.23% 88.02% -1.21% 87.84% 86.59% -1.25% 89.06% 87.89% -1.17% 86.67% 85.33% -1.34% LLaVA-13B 91.82% 90.09% -1.73% 90.72% 90.41% -0.31% 90.10% 91.03% 0.93% 91.35% 89.80% -1.55% MoE-LLaVA 88.46% 86.54% -1.92% 87.16% 86.21% -0.95% 88.28% 86.77% -1.51% 86.07% 85.67% -0.40% MobileVLM 80.13% 80.15% 0.02% 82.88% 82.97% 0.09% 84.14% 83.81% -0.33% 81.67% 82.16% 0.49% Qwen-7B 89.05% 89.06% 0.01% 90.15% 90.57% 0.42% 92.83% 92.77% -0.06% 87.60% 88.46% 0.86% NVILA-8B 99.81% 99.83% 0.02% 99.82% 99.83% 0.01% 99.87% 99.81% -0.06% 99.81% 99.87% 0.06%